# Existence and non-existence of consistent estimators in supercritical controlled branching processes

### v1

@@ 1 Introduction
BGWPs are often not suitable models for biological populations. Indeed, many biological populations do not grow exponentially; for example, due to competition for limited resources, they may exhibit logistic growth. In addition, individuals within the same generation may not give birth independently; for example, this could be due to random population-wide factors, such as weather conditions, that are often referred to as environmental stochasticity [21, Chapter 1.2]. A common extension to the BGWP that overcomes these limitations is the controlled branching process (CBP) $\{Z_{n},z_{0}\}_{n\in\mathbbm{N}_{1}}$, recursively defined as
@@
If $\{\hat{\boldsymbol{\theta}}_{n}\}$ is consistent, then, as more data become available, the sequence converges to the true parameter value $\boldsymbol{\theta}$. On the other hand, if $\{\hat{\boldsymbol{\theta}}_{n}\}$ is not consistent, then we may question whether a consistent estimator for $\boldsymbol{\theta}$ actually exists. If not, this would be an indication that the model may be over-parametrised.
@@
Q1 and Q2 have both been resolved for supercritical BGWPs with and without immigration. Indeed, for a supercritical BGWP without immigration, Lockhart showed (under mild assumptions) that no parameter other than the offspring mean and variance can be estimated consistently [23] (adapted in Theorem 2.1; see also [14, Theorem 1.2]). In addition, Harris [16, Theorem 7.2] and Heyde [17, Theorem 4] showed that the estimator $\hat{m}_{n}:=\sum_{i=1}^{n}Z_{i}/\sum_{i=0}^{n-1}Z_{i}$ is (weakly, resp. strongly) consistent for the offspring mean. A strongly consistent estimator for the offspring variance was established in [18]. For supercritical BGWPs with immigration, Wei and Winnicki [29, Proposition 3.3] showed that no parameter other than the offspring mean and variance can be estimated consistently, thereby extending Lockhart’s result (see also [30, Theorem 4.5] which considers the critical case). In addition, consistent estimators for the offspring mean and variance of these processes were provided in [29, Theorem 2.2] and [18, Section 3], respectively. Consequently, questions about the existence of consistent estimators for supercritical BGWPs with and without immigration have been largely resolved.
@@
Our answers to Q1 help to clarify which parameters might be possible to consistently estimate in each scenario S1S3. In our answers to Q2 we provide explicit expressions for consistent estimators under some additional regularity conditions. In S1 (Theorem 3.5), under minor regularity assumptions, we establish consistent estimators for the mean and variance of the offspring distribution (Theorem 3.5). While the consistency for an estimator of the offspring mean has been previously shown ([10, Theorem 4.2] and [27, Section 6]), albeit under stronger conditions than those we impose, our result for the offspring variance is new: previously, consistency had only been demonstrated in the special case of a deterministic control function [11]. In S2 (Theorem 3.10), under the assumption that $\mathbbm{E}(\phi(z))=\alpha z$ and $\text{Var}(\phi(z))=\beta z$, we establish consistent estimators for the normalised conditional mean and variance of the next step, i.e. for $\mathbb{E}(Z_{1}|Z_{0}=z)/z=m\alpha$ and $\text{Var}(Z_{1}|Z_{0}=z)/z=\sigma^{2}\alpha+m^{2}\beta$, which are the only quantities that can be estimated consistently under the assumptions of Theorem 3.8. In S3 (Theorem 3.13), under similar assumptions as Theorem 3.10, we construct consistent estimators for $m$, $\alpha$, $\sigma^{2}$, and $\beta$: the only quantities that can be estimated consistently under the assumptions of Theorem 3.12. These are the first estimators proven to be consistent under this observation scheme.
@@
Our results have implications in population ecology. In this field, a common rule of thumb is that demographic stochasticity — the randomness inherent in the independent reproduction and lifetime of individuals within a population — and environmental stochasticity — the random changes in environmental conditions that impact a population as a whole — should not be estimated simultaneously from a single trajectory of population size counts [21, Chapter 1.7.1]. To the best of the authors’ knowledge, the justification for this rule has only been empirical. In practice, ecologists use different observation schemes when estimating both types of stochasticity. For example, in studying a bird population, they might estimate demographic stochasticity by counting the clutch size and then treat these demographic parameters as known when estimating environmental stochasticity using population size counts. In the context of CBPs, this principle translates to the idea that both the parameters of $\xi$ (demographic stochasticity) and those of $\phi(\cdot)$ (environmental stochasticity) should not be estimated simultaneously from a single trajectory of population sizes. Our results provide theoretical support for this principle for supercritical CBPs (Q1 for S2), and show that the parameters of $\xi$ and $\phi(\cdot)$ can only be consistently estimated together under a more detailed observation scheme (Q2 for S3), similar to some observation schemes used by ecologists. We believe our arguments can be extended to other stochastic population models such as diffusion models [21] and supercritical branching processes in a random environment [20].
@@
The paper is organised as follows. In Section 2, we outline the fundamental consistency results for supercritical BGWPs and illustrate them with an example. In Section 3.1 we present a general framework for establishing the non-existence of consistent estimators. In Sections 3.2–3.4 we present our answers to Questions Q1 and Q2 for scenarios S1S3. In Section 4, we discuss future work and open questions. Finally, in Section 5 we gather the proofs of our results.
@@ 2.1 Modelling supercritical populations with BGWPs: Whooping cranes
Figure 1: Female population sizes of the whooping crane Aransas-Wood Buffalo flock, 1938–2008 (see [3]),
@@
The next step sizes are convolutions of independent random variables, whose generating functions can therefore be computed easily, and then inverted for example using the numerical technique of Abate and Whitt [1]. By maximising the resulting (approximate) likelihoods, we obtain maximum likelihood estimates (MLEs) for each model:
@@
We use parametric bootstrap [28, Section 13.3] to obtain 95% confidence intervals:
@@
Figure 2: Mean squared error (MSE) of the maximum likelihood estimates (MLEs) generated using parametric bootstrap for Models (i) and (ii) with 1000 simulations of trajectories of length 20, 30, …, 70.
For a supercritical BGWP, it has been shown in [17] and [18], respectively, that consistent estimators for the offspring mean $m$ and offspring variance $\sigma^{2}$ exist. Theorem 2.1 below, adapted from [23, Theorem 2], demonstrates that $m$ and $\sigma^{2}$ are the only parameters of a supercritical BGWP that can be consistently estimated. For Model (i), we can formulate consistent estimators for $p_{0}$, $p_{1}$, and $p_{2}$ in terms of those for $m$ and $\sigma^{2}$, by solving a system of three equations with three unknowns ($\hat{p}_{1}+2\hat{p}_{2}=\hat{m}$,  $\hat{p}_{1}+4\hat{p}_{2}=\hat{\sigma}^{2}+\hat{m}^{2}$,  and $\hat{p}_{0}+\hat{p}_{1}+\hat{p}_{2}=1$). This provides theoretical justification for why the MSE of the estimates in Model (i) appears to converge to zero in Figure 2. For Model (ii), however, we aim to estimate $p_{0}$, $p_{1}$, $p_{2}$, and $p_{3}$ consistently, that is, we now have four unknowns, but we still have only three equations in our system. Theorem 2.1 will then demonstrate that $p_{0}$, $p_{1}$, $p_{2}$, and $p_{3}$ cannot all be consistently estimated simultaneously.
@@
Let us return to Model (ii) in our whooping crane example, with $\boldsymbol{\theta}:=(p_{0},\,p_{1},\,p_{2},\,p_{3})$. Note that if $\{X_{n},z_{0}\}$ is the BGWP with $p_{0,X}=0.1538$, $p_{1,X}=0.6491$, $p_{2,X}=0.1971$ and $p_{3,X}=0$ (matching the MLEs found above), and $\{Z_{n},z_{0}\}$ is a BGWP with $p_{0,Z}=0.0891$, $p_{1,Z}=0.8432$, $p_{2,Z}=0.003$ and $p_{3,Z}=0.0647$, then both processes have the same mean and variance for their offspring distributions. Then, by Theorem 2.1, it is not possible to consistently estimate $\boldsymbol{\theta}$. This provides a theoretical justification for why the MSE of the estimates in Model (ii) appears not to converge to zero in Figure 2.
To understand the intuition behind Theorem 2.1, we note that the observations $(z_{0},\,z_{1},\,\dots,\,z_{70})$ are not taken from the distribution of $\xi$ itself. Instead, they are taken from the distribution of the next-step size $(Z_{n}|Z_{n-1}=z_{n-1})$, which corresponds to the convolution $\sum_{i=1}^{z_{n-1}}\xi_{n,i}$. Given that the population size is growing on $\{Z_{n}\to\infty\}$, and the next-step size distribution is the sum of independent copies of $\xi$, the central limit theorem applies, and therefore all information but the mean and variance of $\xi$ is eventually hidden.
@@ 3.1 A framework for proving the non-existence of consistent estimators
we denote the offspring mean and variance by $m:=\mathbbm{E}\xi$ and $\sigma^{2}:=\text{Var}(\xi)$, assuming throughout that $\sigma^{2}>0$, and we denote the mean and variance of the control function by $\varepsilon(z):=\mathbbm{E}\phi(z)$ and $\nu^{2}(z):=\text{Var}(\phi(z))$. Following [8, p.76], we define the mean growth rate of the process $\{Z_{n},z_{0}\}$ at population size $z$ as
$\tau(z)$ $:=\frac{1}{z}\mathbbm{E}(Z_{1}|Z_{0}=z_{0})$
$=\frac{\varepsilon(z)}{z}\cdot m,$
@@
Recall that $\{Z_{n},z_{0}\}$ is said to grow unboundedly if $Z_{n}\to\infty$ as $n\to\infty$. Unlike BGWPs, supercritical CBPs do not necessarily have a positive probability of unbounded growth (see [8, Example 3.1]). Theorem 3.2 of [8] provides a sufficient condition for $\mathbbm{P}(Z_{n}\to\infty)>0$, namely that there exist $a,b>0$ such that
@@
In fact, [8, Theorem 3.2] shows that under (6), $\mathbbm{P}(Z_{n}\to\infty)\to 1$ as the initial population size $z_{0}$ approaches infinity.
@@
We now relate the total variation distance between the distributions of two processes $\{Z_{n},z_{0}\},\{X_{n},z_{0}\}\in\Pi$,
{$||\mathcal{L}_{\{Z_{n},z_{0}\}}-\mathcal{L}_{\{X_{n},z_{0}\}}||_{TV}:=\sup_{C\subseteq\mathbbm{N}_{0}^{\infty}}|\mathbbm{P}(\{Z_{n},z_{0}\}\in C)-\mathbbm{P}(\{X_{n},z_{0}\}\in C)|,$
to the non-existence of consistent estimators for $\boldsymbol{\theta}$. If a consistent estimator $\hat{\boldsymbol{\theta}}_{n}$ exists, then we can solve the following simple classification problem: Given an infinite trajectory generated from either $\{Z_{n},z_{0}\}$ or $\{X_{n},z_{0}\}$ with $\boldsymbol{\theta}_{Z}\neq\boldsymbol{\theta}_{X}$, can we identify which process generated the trajectory with arbitrarily high accuracy? If a consistent estimator exists, then the answer is positive. This is because if $\hat{\boldsymbol{\theta}}_{n}$ converges to $\boldsymbol{\theta}_{Z}$ (resp. to $\boldsymbol{\theta}_{X}$), then we know that the trajectory was generated by $\{Z_{n},z_{0}\}$ (resp. by $\{X_{n},z_{0}\}$). However, if {$||\mathcal{L}_{\{Z_{n},z_{0}\}}-\mathcal{L}_{\{X_{n},z_{0}\}}||_{TV}<1$, then it is not possible to always make the correct classification. This can be seen through an analogy with the univariate setting: given an observation $x\in\mathbb{R}$, we want to determine whether this observation was generated from either $X_{1}\sim f_{\theta_{1}}(x)$ or $X_{2}\sim f_{\theta_{2}}(x)$. If $||\mathcal{L}_{X_{1}}-\mathcal{L}_{X_{2}}||_{TV}<1$, as on the left-hand-side of Figure 4, it is not always possible to correctly classify the observation, whereas if $||\mathcal{L}_{X_{1}}-\mathcal{L}_{X_{2}}||_{TV}=1$, as on the right-hand-side of Figure 4, it is always possible. Coming back to CBPs, if we are not able to classify an infinitely long trajectory as coming from $\{Z_{n},z_{0}\}$ or $\{X_{n},z_{0}\}$ (i.e. because {$||\mathcal{L}_{\{Z_{n},z_{0}\}}-\mathcal{L}_{\{X_{n},z_{0}\}}||_{TV}<1$), then no consistent estimator exists for $\boldsymbol{\theta}$.
$x$$f_{\theta_{1}}(x)$$f_{\theta_{2}}(x)$
$x$$f_{\theta_{1}}(x)$$f_{\theta_{2}}(x)$
Figure 4: Left: Two overlapping densities with $||\mathcal{L}_{X_{1}}-\mathcal{L}_{X_{2}}||_{TV}<1$; we do not know if the observation $x$ was generated from $f_{\theta_{1}}$ or $f_{\theta_{2}}$. Right: Two non-overlapping densities with $||\mathcal{L}_{X_{1}}-\mathcal{L}_{X_{2}}||_{TV}=1$; we can be certain that the observation $x$ was generated from $f_{\theta_{1}}$.
@@
$||\mathcal{L}_{Z_{1}|Z_{0}=z_{0}}-\mathcal{L}_{X_{1}|X_{0}=z_{0}}||_{TV}=O(z_{0}^{-q})\quad\textrm{for some $q>0$.}$ (7)
Then
$\lim_{z_{0}\to\infty}||\mathcal{L}_{\{Z_{n},z_{0}\}}-\mathcal{L}_{\{X_{n},z_{0}\}}||_{TV}=0.$
In the context of our discussion above, note that if $\lim_{z_{0}\to\infty}||\mathcal{L}_{\{Z_{n},z_{0}\}}-\mathcal{L}_{\{X_{n},z_{0}\}}||_{TV}=0$, then {$||\mathcal{L}_{\{Z_{n},z_{0}\}}-\mathcal{L}_{\{X_{n},z_{0}\}}||_{TV}<1$ for all sufficiently large $z_{0}$. By combining Lemma 3.1 with the above relationship between the total variation distance and the non-existence of consistent estimators, we obtain the next result, which we will further refine in Sections 3.23.4 (see Theorems 3.4, 3.8, and 3.12).

### v2

@@ 1 Introduction
BGWPs are often not suitable models for biological populations. Indeed, many biological populations do not grow exponentially; for example, due to competition for limited resources, they may exhibit logistic growth. In addition, individuals within the same generation may not give birth independently; for example, this could be due to random population-wide factors, such as weather conditions, that are often referred to as environmental stochasticity [23, Chapter 1.2]. A common extension to the BGWP that overcomes these limitations is the controlled branching process (CBP). Using the definition introduced by Yanev [33] (see also [29]), a CBP $\{Z_{n},z_{0}\}_{n\in\mathbbm{N}_{1}}$ is defined recursively as
@@
If $\{\hat{\boldsymbol{\theta}}_{n}\}$ is consistent, then, as more data become available, the sequence converges to the true parameter value $\boldsymbol{\theta}$. On the other hand, if $\{\hat{\boldsymbol{\theta}}_{n}\}$ is not consistent, then we may question whether a consistent estimator for $\boldsymbol{\theta}$ actually exists. If not, this may be an indication that the model is over-parametrised.
@@
Q1 and Q2 have both been resolved for supercritical BGWPs with and without immigration. Indeed, for a supercritical BGWP without immigration, Lockhart showed (under mild assumptions) that no parameter other than the offspring mean and variance can be estimated consistently [25] (adapted in Theorem 2.1; see also [15, Theorem 1.2]). In addition, Harris [18, Theorem 7.2] and Heyde [19, Theorem 4] showed that the estimator $\hat{m}_{n}:=\sum_{i=1}^{n}Z_{i}/\sum_{i=0}^{n-1}Z_{i}$ is (weakly, resp. strongly) consistent for the offspring mean. A strongly consistent estimator for the offspring variance was established in [20]. For supercritical BGWPs with immigration, Wei and Winnicki [31, Proposition 3.3] showed that no parameter other than the offspring mean and variance can be estimated consistently, thereby extending Lockhart’s result (see also [32, Theorem 4.5] which considers the critical case). In addition, consistent estimators for the offspring mean and variance of these processes were provided in [31, Theorem 2.2] and [20, Section 3], respectively. Consequently, questions about the existence of consistent estimators for supercritical BGWPs with and without immigration have been largely resolved.
@@
Our answers to Q1 help to clarify which parameters might be possible to consistently estimate in each scenario S1S3. In our answers to Q2 we provide explicit expressions for consistent estimators under some additional regularity conditions. In S1 (Theorem 3.5), under minor regularity assumptions, we establish consistent estimators for the mean and variance of the offspring distribution (Theorem 3.5). Consistent estimators for the offspring mean have been derived in [11, Theorem 4.2] and [30, Section 6]; however, both assume that the limit $\tau:=\lim_{z\to\infty}\mathbbm{E}(Z_{1}|Z_{0}=z)/z$ exists, and that $\lim_{n\to\infty}(\tau m)^{-n}Z_{n}$ converges to a non-degenerate random variable. Our consistent estimator for the mean holds without these restrictive assumptions. For the offspring variance, consistency had only been demonstrated in the special case of a deterministic control function [12]. Our consistent estimator for the variance holds for a random control function. Our results require a different proof approach than what has previously been used in the literature. In S2 (Theorem 3.10), under the assumption that $\mathbbm{E}(\phi(z))=\alpha z$ and $\text{Var}(\phi(z))=\beta z$, we establish consistent estimators for the normalised conditional mean and variance of the next step, i.e. for $\mathbb{E}(Z_{1}|Z_{0}=z)/z=m\alpha$ and $\text{Var}(Z_{1}|Z_{0}=z)/z=\sigma^{2}\alpha+m^{2}\beta$, which are the only quantities that can be estimated consistently under the assumptions of Theorem 3.8. In S3 (Theorem 3.13), under similar assumptions as Theorem 3.10, we construct consistent estimators for $m$, $\alpha$, $\sigma^{2}$, and $\beta$: the only quantities that can be estimated consistently under the assumptions of Theorem 3.12. These are the first estimators proven to be consistent under this observation scheme.
@@
Our results have implications in population ecology. In this field, a common rule of thumb is that demographic stochasticity—the randomness inherent in the independent reproduction and lifetime of individuals within a population—and environmental stochasticity—the random changes in environmental conditions that impact a population as a whole—should not be estimated simultaneously from a single trajectory of population size counts [23, Chapter 1.7.1]. To the best of the authors’ knowledge, the justification for this rule has only been empirical. In practice, ecologists use different observation schemes when estimating both types of stochasticity. For example, in studying a bird population, they might estimate demographic stochasticity by counting the clutch size and then treat these demographic parameters as known when estimating environmental stochasticity using population size counts. In the context of CBPs, this principle translates to the idea that both the parameters of $\xi$ (demographic stochasticity) and those of $\phi(\cdot)$ (environmental stochasticity) should not be estimated simultaneously from a single trajectory of population sizes. Our results provide theoretical support for this principle for supercritical CBPs (Q1 for S2), and show that the parameters of $\xi$ and $\phi(\cdot)$ can only be consistently estimated together under a more detailed observation scheme (Q2 for S3), similar to some observation schemes used by ecologists. We believe our arguments can be extended to other stochastic population models such as diffusion models [23] and supercritical branching processes in a random environment [22].
@@
The paper is organised as follows. In Section 2, we outline the fundamental consistency results for supercritical BGWPs and illustrate them with an example. In Section 3.1 we present a general framework for establishing the non-existence of consistent estimators. In Sections 3.2–3.4 we present our answers to Questions Q1 and Q2 for scenarios S1S3. In Section 4, we discuss future work and open questions. In Section 5 we gather the proofs for each of our non-existence results, while proofs for the consistency of estimators can be found in Section 6.
@@ 2.1 Modelling supercritical populations with BGWPs: Whooping cranes
Figure 1: Female population sizes of the whooping crane Aransas-Wood Buffalo flock, 1938–2008 (see [3]).
@@
These next step sizes are convolutions of independent random variables, whose generating functions can therefore be computed easily, and then inverted for example using the numerical technique of Abate and Whitt [1]. By maximising the resulting (approximate) likelihoods, we obtain maximum likelihood estimates (MLEs) for each model:
@@
We use parametric bootstrap [7, Section 13.3] to obtain 95% confidence intervals:
@@
Figure 2: Mean squared error of the maximum likelihood estimates generated using parametric bootstrap for Models (i) and (ii) with 1000 simulations of trajectories of length 20, 30, …, 70.
For a supercritical BGWP, it has been shown in [19] and [20], respectively, that consistent estimators for the offspring mean $m$ and offspring variance $\sigma^{2}$ exist. Theorem 2.1 below, adapted from [25, Theorem 2], demonstrates that $m$ and $\sigma^{2}$ are the only parameters of a supercritical BGWP that can be consistently estimated. For Model (i), we can formulate consistent estimators for $p_{0}$, $p_{1}$, and $p_{2}$ in terms of those for $m$ and $\sigma^{2}$, by solving a system of three equations with three unknowns ($\hat{p}_{1}+2\hat{p}_{2}=\hat{m}$,  $\hat{p}_{1}+4\hat{p}_{2}=\hat{\sigma}^{2}+\hat{m}^{2}$,  and $\hat{p}_{0}+\hat{p}_{1}+\hat{p}_{2}=1$). This provides theoretical justification for why the MSE of the estimates in Model (i) appears to converge to zero in Figure 2. For Model (ii), we aim to estimate $p_{0}$, $p_{1}$, $p_{2}$, and $p_{3}$ consistently using $\hat{m}$ and $\hat{\sigma}^{2}$, however, we now have four unknowns but we still have only three equations in our system. Theorem 2.1 will then demonstrate that $p_{0}$, $p_{1}$, $p_{2}$, and $p_{3}$ cannot all be consistently estimated simultaneously.
@@
Let us return to Model (ii) in the whooping crane example, with $\boldsymbol{\theta}:=(p_{0},\,p_{1},\,p_{2},\,p_{3})$. Note that if $\{X_{n},z_{0}\}$ is the BGWP with $p_{0,X}=0.1538$, $p_{1,X}=0.6491$, $p_{2,X}=0.1971$ and $p_{3,X}=0$ (matching the MLEs found above), and $\{Z_{n},z_{0}\}$ is a BGWP with $p_{0,Z}=0.0891$, $p_{1,Z}=0.8432$, $p_{2,Z}=0.003$ and $p_{3,Z}=0.0647$, then both processes have the same mean and variance for their offspring distributions. Thus, by Theorem 2.1, it is not possible to consistently estimate $\boldsymbol{\theta}$. This provides a theoretical justification for why the MSE of the estimates in Model (ii) appears not to converge to zero in Figure 2.
To understand the intuition behind Theorem 2.1, we note that the observations $(z_{0},\,z_{1},\,\dots,\,z_{70})$ are not taken from the distribution of $\xi$ itself. Instead, they are taken from the distribution of the next-step size $(Z_{n}|Z_{n-1}=z_{n-1})$, which corresponds to the convolution $\sum_{i=1}^{z_{n-1}}\xi_{n,i}$. Given that the population size is growing on $\{Z_{n}\to\infty\}$, and the next-step size distribution is the sum of independent copies of $\xi$, the central limit theorem applies, and thus all information but the mean and variance of $\xi$ is eventually hidden.
@@ 3.1 A framework for proving the non-existence of consistent estimators
we denote the offspring mean and variance by $m:=\mathbbm{E}\xi$ and $\sigma^{2}:=\text{Var}(\xi)$, assuming throughout that $\sigma^{2}>0$, and we denote the mean and variance of the control function by $\varepsilon(z):=\mathbbm{E}\phi(z)$ and $\nu^{2}(z):=\text{Var}(\phi(z))$. Following [9, p.76], we define the mean growth rate of the process $\{Z_{n},z_{0}\}$ at population size $z$ as
$\tau(z):=\frac{1}{z}\mathbbm{E}(Z_{1}|Z_{0}=z)=\frac{\varepsilon(z)}{z}\cdot m,$
@@
Recall that $\{Z_{n},z_{0}\}$ is said to grow unboundedly if $Z_{n}\to\infty$ as $n\to\infty$. Unlike BGWPs, supercritical CBPs do not necessarily have a positive probability of unbounded growth (see [9, Example 3.1]). Theorem 3.2 of [9] provides a sufficient condition for $\mathbbm{P}(Z_{n}\to\infty)>0$, namely that there exist $a,b>0$ such that
@@
In fact, [9, Theorem 3.2] shows that under (6), $\mathbbm{P}(Z_{n}\to\infty)\to 1$ as the initial population size $z_{0}$ approaches infinity.
@@
We now relate the total variation distance (TVD) between the distributions of two processes $\{Z_{n},z_{0}\}$, $\{X_{n},z_{0}\}\in\Pi$,
$\mathcal{L}_{\{Z_{n},z_{0}\}}-\mathcal{L}_{\{X_{n},z_{0}\}}||_{TV}:=\sup_{C\subseteq\mathbbm{N}_{0}^{\infty}}|\mathbbm{P}(\{Z_{n},z_{0}\}\in C)-\mathbbm{P}(\{X_{n},z_{0}\}\in C)|,$
to the non-existence of consistent estimators for $\boldsymbol{\theta}$. If a consistent estimator $\hat{\boldsymbol{\theta}}_{n}$ exists, then we can solve the following simple classification problem: Given an infinite trajectory generated from either $\{Z_{n},z_{0}\}$ or $\{X_{n},z_{0}\}$ with $\boldsymbol{\theta}_{Z}\neq\boldsymbol{\theta}_{X}$, can we identify which process generated the trajectory with arbitrarily high accuracy? If a consistent estimator exists, then the answer is positive. This is because if $\hat{\boldsymbol{\theta}}_{n}$ converges to $\boldsymbol{\theta}_{Z}$ (resp. to $\boldsymbol{\theta}_{X}$), then we know that the trajectory was generated by $\{Z_{n},z_{0}\}$ (resp. by $\{X_{n},z_{0}\}$). However, if $\mathcal{L}_{\{Z_{n},z_{0}\}}-\mathcal{L}_{\{X_{n},z_{0}\}}||_{TV}<1$, then it is not possible to always make the correct classification. This can be seen through an analogy with the univariate setting: given an observation $x\in\mathbb{R}$, we want to determine whether this observation was generated from either $X_{1}\sim f_{\theta_{1}}(x)$ or $X_{2}\sim f_{\theta_{2}}(x)$. If $\mathcal{L}_{X_{1}}-\mathcal{L}_{X_{2}}||_{TV}<1$, as on the left-hand-side of Figure 4, it is not always possible to correctly classify the observation, whereas if $\mathcal{L}_{X_{1}}-\mathcal{L}_{X_{2}}||_{TV}=1$, as on the right-hand-side of Figure 4, it is always possible. Coming back to CBPs, if we are not able to classify an infinitely long trajectory as coming from $\{Z_{n},z_{0}\}$ or $\{X_{n},z_{0}\}$ (i.e. because $\mathcal{L}_{\{Z_{n},z_{0}\}}-\mathcal{L}_{\{X_{n},z_{0}\}}||_{TV}<1$), then no consistent estimator exists for $\boldsymbol{\theta}$.
[…]
Figure 4: Left: Two overlapping densities with $\mathcal{L}_{X_{1}}-\mathcal{L}_{X_{2}}||_{TV}<1$; we do not know if the observation $x$ was generated from $f_{\theta_{1}}$ or $f_{\theta_{2}}$. Right: Two non-overlapping densities with $\mathcal{L}_{X_{1}}-\mathcal{L}_{X_{2}}||_{TV}=1$; we can be certain that the observation $x$ was generated from $f_{\theta_{1}}$.
@@
$\mathcal{L}_{Z_{1}|Z_{0}=z_{0}}-\mathcal{L}_{X_{1}|X_{0}=z_{0}}||_{TV}=O(z_{0}^{-q})\quad\textrm{for some $q>0$.}$ (7)
[…]